OpenStudy (anonymous):

Find (g o g)(x) for the function g(x)= 4/x+7

myininaya (myininaya):

Is that \[g(x)=\frac{4}{x+7} \text{ or } g(x)=\frac{4}{x}+7 \]

OpenStudy (anonymous):

The first! It's (x+7) Sorry about the confusion!

myininaya (myininaya):

\[g(x)=\frac{4}{x+7}\] \[g(g(x))=g(\frac{4}{x+7})\]

myininaya (myininaya):

Now where ever theere is an x in g replace it with 4/(x+7)

myininaya (myininaya):

Like.... \[g(g(x))=\frac{4}{\frac{4}{x+7}+7}\]

myininaya (myininaya):

multiply top and bottom by (x+7) to fix that fraction up a bit

OpenStudy (anonymous):

so the answer is 4(x+7)/(7x+53) ?

myininaya (myininaya):

\[\frac{4(x+7)}{4+7(x+7)}\] Yes thats right! great job! :)